Last year, 20 percent of all teenagers purchased a new iPhone. This year, a sample of 260 randomly chosen teenagers showed that 39 had purchased a new iPhone. To test whether the percentage has decreased, the p-value is:
Answer #1

We need to test if the percentage has decreased. Hence, our Hypotheses will be:

Ho: p = 0.20

Ha: p < 0.20

Sample proportion = 39/260 = 0.15

Test statistic

z = -2.02

Hence,

p-Value = P(z < -2.02) = 0.0219